A local building code requires that all schools must be surrounded by a lawn. The width of the lawn must
be uniform and the area of the lawn must equal the area of the school. What must be the width of a lawn
surrounding a rectangular school that measures 120 m by 80 m?

Answer by Boreal(15235)

Drawing this helps.
The area of the school is 9600 m^2.
The area of the lawn must be the same, so the total area must be 19200 m^2.
The whole 19200 ft^2 is 120+2x by 80+2x, where x equals the uniform width of the lawn.
That product is 9600+400x+4x^2=19200
4x^2+400-9600=0
divide by 4
x^2+100-2400=0
(x+120)(x-20)=0
take positive root only. x=20 m ANSWER
The width of the whole area is 160 x 120, and that is 19200 m^2.